OpenStudy (happpily):

A candy store makes a 9-pound mixture of gummy candy, jelly beans, and hard candy. The cost of gummy candy is $2.00 per pound, jelly beans cost $3.00 per pound, and hard candy costs $3.00 per pound. The mixture calls for two times as many gummy candy pieces as jelly beans. The total cost of the mixture is $23.00. How much of each ingredient did the store use?

OpenStudy (anonymous):

9-pound mixture of gummy candy, jelly beans, and hard candy. Gummy candy = x jelly beans = y hard candy = z x+y +z = 9 --------(1) $2.00 per pound, jelly beans cost $3.00 per pound, and hard candy costs $3.00 per pound 2x +3y +3z = 23. -------- (2) The mixture calls for two times as many gummy candy pieces as jelly beans. y = 2x ------- (3) Putting (3) in (2) and (1) 2x +x+z = 9 3x +z = 9 ------ (4) 2x +3(2x) +3z =23 2x +6x +3z = 23 8x + 3z = 23 ----- (5) simplify (4) and (5) then use the value of x in (3) to find y

OpenStudy (happpily):

ohhhhhhh okay thank you.